# Large Language Models Are Struggle to Cope with Unreasonability in Math Problems

### v5

@@ 5.1 Probing Study
Before exploring how to improve model behavior on unreasonable problems, we first ask a fundamental question: do models even understand what counts as unreasonable? If a model lacks this awareness, it is unlikely to respond appropriately when encountering flawed questions. To test this, we design a simple probing task, where the model is presented with short, isolated expressions that are syntactically valid but semantically unreasonable—for example, “the voting result is -20 votes” or “there are 2.5 people.” We refer to this metric as Unreasonable Phrase Accuracy (UPA). For each expression, the model is asked to judge whether it makes sense.
We construct this test set by extracting unreasonable phrases from the UMP benchmark and evaluate a representative subset of models used in our main experiments (Section 3), including general-purpose, reasoning-enhanced, and math-specialized models. As shown in Figure 5, general-purpose and reasoning-enhanced models typically perform well on this task, suggesting that they retain basic commonsense priors. In contrast, several math-specialized models perform noticeably worse, indicating that domain-specific fine-tuning may come at the cost of general semantic awareness.
@@ 5.3 Fine-Tuning Based Method
We further investigate whether the model’s ability to detect irrationality can be enhanced through supervised fine-tuning. To construct the training data, we pair unreasonable questions from GSM8K with carefully edited responses based on DeepSeek-R1 outputs. These responses explicitly identify the flaws through concise, step-by-step reasoning, encouraging appropriate reflection without triggering overthinking. We also include original questions paired with similarly edited answers to maintain general reasoning competence and prevent the model from overgeneralizing flaw detection to well-posed inputs. This balanced training setup provides the model with positive demonstrations of how to respond to both flawed and valid inputs. We fine-tune on a subset of GSM8K and evaluate the resulting model on MATH to assess generalization beyond the training distribution. As shown in Figure 6, the fine-tuned model exhibits clear improvement in detecting unreasonable inputs. However, this gain comes at a cost: accuracy on original problems drops noticeably. This suggests a trade-off between flaw sensitivity and general problem-solving ability. Further training configurations and examples are detailed in Appendix F.
@@ Discussion.
6 Related Work
6.1 LLMs as Math Problem Solvers
@@
6.2 Improve Models’ Inference Ability
@@
6.3 Model Refusal Phenomenon
The phenomenon of model refusal, where LLMs decline to answer certain prompts is central to ensuring AI safety and reliability. Refusals may stem from safety training, uncertainty or knowledge gaps or defects in the prompt itself. Recent studies have explored both linear and nonlinear mechanisms behind refusal behaviors, including "refusal directions" in activation space and the use of "refusal tokens" to control response behavior at inference. However, over-refusal—where benign prompts are incorrectly rejected remains a key challengePasch (2025); Hildebrandt et al. (2025). Approaches like RAIT, CRaFT, and Think-Before-Refusal aim to reduce false refusals by incorporating reasoning and certainty estimationZhu et al. (2024); Zhang et al. (2024a). Evaluation efforts have led to dedicated benchmarks such as OR-Bench, SORRY-Bench, and MM-UPD, which assess refusal quality under various conditions, highlighting the need for context-aware and nuanced assessment frameworksCui et al. (2024); Xie et al. (2025); Miyai et al. (2025).
7 Conclusion
In this work, we investigate how large language models behave when confronted with unreasonable mathematical problems. To facilitate this, we construct a benchmark comprising over 1,000 questions containing hidden irrationality. Experimental results reveal that most models struggle to identify such unreasonableness. While reasoning-enhanced models are more likely to uncover hidden inconsistencies during multi-step reasoning, they often fall into repetitive reflection and overthinking behavior that ultimately hinders clarity and usefulness. To further explore whether models possess latent capabilities for detecting flaws, we experiment with critical-thinking prompts and supervised fine-tuning. Our findings suggest that many models do have the potential to detect unreasonable content, but this ability requires explicit activation. We hope that our benchmark serves as a valuable tool for evaluating both the reliability and behavioral robustness of LLMs in the presence of unreasonable inputs.
Limitations
This study has several limitations. First, we do not include proprietary reasoning models such as those from OpenAI in our analysis of reasoning-enhanced systems. Since these APIs do not expose intermediate reasoning chains, comparing their behavior to open-source models that explicitly generate multi-step thoughts would introduce unfairness in both behavioral and metric-based evaluations. Second, due to hardware constraints, our intervention experiments such as prompt engineering and supervised fine-tuning are conducted only on 7B-scale models. While these experiments provide encouraging insights, scaling these methods to larger models and exploring more effective solutions for mitigating failure behaviors remain important directions for future work.
Ethical Considerations
Our paper explores how llms perform in the face of unreasonable mathematical problems. Sometimes the model may not realize the irrationality in the math problem, so it may produce wrong answers and mislead the user.
Appendix A Category of UMP Benchmark
Below is the description of the five categories. Each of our data pieces is made up of five parts. An original question paired with its corresponding answer , along with a designated type of unreasonableness we aim to investigate, we systematically construct a rephrased, unreasonable question and provide an accompanying explanation . Consequently, our benchmark comprises quintuples of the form .
Appendix B Probabilistic View from Prior Conditioning
@@
Appendix C Model Versions and Inference Settings
We include both API-access and open-source models in our evaluation. The API versions used are gpt-4o-2024-11-20, grok-3-reasoning and deepseek-r1-2025-01-20, claude-3.5-sonnet-20241022. To maximize accuracy on original (well-posed) questions, we adopt temperature = 0.0(greedy decoding) for all open-source non-reasoning models. However, for reasoning-enhanced models, we observe that greedy decoding may result in unstable behaviors, such as excessively long outputs, especially in unreasonable problems. Based on these findings, we set temperature = 0.6 for all reasoning models during evaluation.(We provide further analysis in Appendix H.) We set , which means that the model must get all three responses right. For API-based models, we maintain their default decoding settings. Due to access limitations and cost constraints, all the responses are generated in a single pass without sampling-based aggregation.
Appendix D Keywords for Reflection Detection
@@
Appendix E Similarity between Original Problems and Unreasonable Problems
@@
Appendix F Experiment Setup for Supervised Fine-Tuning
@@
Appendix G Template for Critical Prompting
@@
Appendix H Effect of Temperature on Model Behavior
@@
Appendix I Case Study
I.1 An Example of Overthinking
@@
I.2 The unstable of math-specified model
@@
I.3 Model Unconsciousness of Problem Errors
Tables 9, 10, present examples where the model completely fails to recognize the errors embedded in the questions. These cases suggest that the model entirely overlooks the unreasonableness of the input and proceeds as if the problems were well-posed.
@@ I.3 Model Unconsciousness of Problem Errors
Appendix J Example output by different LLM
In this section, we will demonstrate how various models perform on unreasonable problems using the conventional Chain of Thought approach compared to adding a critical instruction method. Tables 16 display the outputs from Qwen2.5-7B-Instruct using two different prompts. It is evident that the outputs generated by adding critical instruction exhibit marked enhancements.
@@
Appendix K Evaluation Prompt Details
@@
Appendix L Prompt for Unreasonable Problem Generation
@@
Appendix M Examples in UMP benchmark
@@
Appendix N Query format for Pre-validation experiment

### v6

@@ 5.1 Probing Study
Before exploring how to improve model behavior on unreasonable problems, we first ask a fundamental question: do models even understand what counts as unreasonable? If a model lacks this awareness, it is unlikely to respond appropriately when encountering flawed questions. To test this, we design a simple probing task, where the model is presented with short, isolated expressions that are syntactically valid but semantically unreasonable—for example, “the voting result is -20 votes” or “there are 2.5 people.” We refer to this metric as Unreasonable Phrase Accuracy (UPA). For each expression, the model is asked to judge whether it makes sense. We construct this test set by extracting unreasonable phrases from the UMP benchmark and evaluate a representative subset of models used in our main experiments (Section 3), including general-purpose, reasoning-enhanced, and math-specialized models. As shown in Figure 5, general-purpose and reasoning-enhanced models typically perform well on this task, suggesting that they retain basic commonsense priors. In contrast, several math-specialized models perform noticeably worse, indicating that domain-specific fine-tuning may come at the cost of general semantic awareness.
@@ 5.3 Fine-Tuning Based Method
We further investigate whether the model’s ability to detect unreasonability can be enhanced through supervised fine-tuning. To construct the training data, we pair unreasonable questions from GSM8K with carefully edited responses based on DeepSeek-R1 outputs. These responses explicitly identify the flaws through concise, step-by-step reasoning, encouraging appropriate reflection without triggering overthinking. We also include original questions paired with similarly edited answers to maintain general reasoning competence and prevent the model from overgeneralizing flaw detection to well-posed inputs. This balanced training setup provides the model with positive demonstrations of how to respond to both flawed and valid inputs. We fine-tune on a subset of GSM8K and evaluate the resulting model on MATH to assess generalization beyond the training distribution. As shown in Figure 6, the fine-tuned model exhibits clear improvement in detecting unreasonable inputs. However, this gain comes at a cost: accuracy on original problems drops noticeably. This suggests a trade-off between flaw sensitivity and general problem-solving ability. Further training details are shown in Appendix I.
@@ Discussion.
6 Conclusion
In this work, we investigate how LLMs behave when confronted with unreasonable mathematical problems. To facilitate this, we construct a benchmark comprising over 1,000 questions containing hidden unreasonability. Experimental results reveal that most models struggle to identify such unreasonableness. While reasoning-enhanced models are more likely to uncover hidden inconsistencies during multi-step reasoning, they often fall into repetitive reflection and overthinking behavior that ultimately hinders clarity and usefulness. To further explore whether models possess latent capabilities for detecting flaws, we experiment with critical-thinking prompts and supervised fine-tuning. Our findings suggest that many models do have the potential to detect unreasonable content, but this ability requires explicit activation. We hope that our benchmark serves as a valuable tool for evaluating both the trustworthiness and behavioral robustness of LLMs in the presence of unreasonable inputs.
Limitations
This study has several limitations. First, we do not include proprietary reasoning models such as those from OpenAI in our analysis of reasoning-enhanced systems. Since these APIs do not expose intermediate reasoning chains, comparing their behavior to open-source models that explicitly generate multi-step reasoning would introduce unfairness in both behavioral and metric-based evaluations. Second, due to hardware constraints, our intervention experiments such as prompt engineering and supervised fine-tuning are conducted only on 7B-scale models. While these experiments provide encouraging insights, scaling these methods to larger models and exploring more effective solutions for mitigating failure behaviors remain important directions for future work.
Ethical Considerations
Our paper explores how LLMs perform in the face of unreasonable mathematical problems. Sometimes the model may not realize the unreasonability in the math problem, which may result in incorrect answers that mislead users.
Appendix A Category of UMP Benchmark
Below is the description of the five categories. Each data point consists of an original question paired with its corresponding answer , a designated type of unreasonableness , a rephrased unreasonable version , and an explanation . Consequently, our benchmark comprises quintuples of the form .
Appendix B Related Work
B.1 LLMs as Math Problem Solvers
@@
B.2 Improve Models’ Inference Ability
@@
B.3 Model Refusal Phenomenon
The phenomenon where LLMs decline to answer certain prompts is central to ensuring AI safety and reliability. Refusals may stem from safety training, uncertainty, knowledge gaps or defects in the prompt itself. Recent studies have explored both linear and nonlinear mechanisms behind refusal behaviors, including "refusal directions" in activation space and the use of "refusal tokens" to control response behavior at inference. However, over-refusal-where benign prompts are incorrectly rejected- remains a key challengePasch (2025); Hildebrandt et al. (2025). Approaches like RAIT, CRaFT, and Think-Before-Refusal aim to reduce false refusals by incorporating reasoning and certainty estimationZhu et al. (2024); Zhang et al. (2024a). Evaluation efforts have led to dedicated benchmarks such as OR-Bench, SORRY-Bench, and MM-UPD, which assess refusal quality under various conditions, highlighting the need for context-aware and nuanced assessment frameworksCui et al. (2024); Xie et al. (2025); Miyai et al. (2025).
Appendix C Probabilistic View from Prior Conditioning
@@
Appendix D Model Versions and Inference Settings
We include both API-access and open-source models in our evaluation. The API versions used are gpt-4o-2024-11-20, grok-3-reasoning and deepseek-r1-2025-01-20, claude-3.5-sonnet-20241022. To maximize accuracy on original (well-posed) questions, we adopt temperature = 0.0(greedy decoding) for all open-source non-reasoning models. However, for reasoning-enhanced models, we observe that greedy decoding may result in unstable behaviors, such as excessively long outputs, especially in unreasonable problems. Based on these findings, we set temperature = 0.6 for all reasoning models during evaluation.(We provide further analysis in Appendix K.) We set , which means that the model must get all three responses right. For API-based models, we maintain their default decoding settings. Due to access limitations and cost constraints, all the responses are generated in a single pass without sampling-based aggregation.
Appendix E Does Solving the Original Help?
Appendix F Keywords for Reflection Detection
@@
Appendix G Similarity between Original Problems and Unreasonable Problems
@@
Appendix H Annotation Protocol and Agreement
To ensure the quality of the benchmark, three reviewers with mathematical training participated in a multi-stage validation process. The following criteria were jointly defined and applied during annotation:
1.
Remove examples that remain mathematically reasonable after rewriting but are mistakenly flagged as unreasonable by the model.
2.
Remove examples where the unreasonable elements are too obvious or violate basic commonsense (e.g., “2.33 people”).
3.
Remove examples where the rewritten version differs too much from the original, failing to preserve surface similarity.
4.
Revise examples with unclear or poorly explained unreasonable elements to make the flaw logically coherent and minimally invasive.
An example is retained only if at least two out of three reviewers judged it as valid under the above criteria.
Appendix I Experiment Setup for Supervised Fine-Tuning
@@
Appendix J Template for Critical Prompting
@@
Appendix K Effect of Temperature on Model Behavior
@@
Appendix L Case Study
L.1 Examples of Overthinking
@@
L.2 The Unstability of Math-Specified Model
@@
L.3 Model Unconsciousness of Problem Errors
Tables 9 and 10, present examples where the model completely fails to recognize the errors embedded in the questions. These cases suggest that the model entirely overlooks the unreasonableness of the input and proceeds as if the problems were well-posed.
@@
Appendix M Example output by different LLM
In this section, we demonstrate how various models perform on unreasonable problems using the conventional Chain of Thought approach compared to adding a critical instruction method. Table 16 displays the outputs from Qwen2.5-7B-Instruct using two different prompts. It is evident that the outputs generated by adding critical instruction exhibit marked enhancements.
@@
Appendix N Evaluation Prompt Details
@@
Appendix O Prompt for Unreasonable Problem Generation
@@
Appendix P Examples in UMP benchmark
@@
Appendix Q Query format for Pre-validation experiment